# Acoustic pulse propagation in a non-ideal shallow-water model

### v1

@@ Abstract
Keywords: acoustic wave propagation; shallow-water acoustics; $\varepsilon$-pseudodifferential operators; Hamiltonian systems; non-self-adjoint operators; WKB–Maslov method; dispersion; ray tracing; frequency and geometric dispersion; perturbation theory; semiclassical analysis
@@ Acknowledgments
A. Kaplun is supported by the Center for Integration in Science, Israel Ministry of Aliyah and Integration and the post-doctoral scholarship at the Leon H. Charney School of Marine Sciences, University of Haifa.
B. Katsnelson is supported by ISF-946/20.
Introduction
@@
Classical formulations of underwater sound propagation rely on ray and mode theories [weinberg1974horizontal, burridge2005horizontal, vcerveny2001seismic, vcerveny2020seismic], in which the acoustic field is represented as a superposition of horizontally propagating rays and vertically standing modes. The theoretical foundation for such methods was established in the 1970s–1980s through the development of the space–time ray method by Babich, Molotkov, and Buldyrev [babich1998space, babich1980space], and the horizontal ray formalism of Weinberg and Burridge [weinberg1974horizontal]. These approaches were later generalized to complex and moving boundary problems [vcerveny1982space, kirpichnikova1983reflection, babich1984complex, farra1989ray, farra1993ray] and further extended into hybrid ray–mode formulations for inhomogeneous ducts [felsen1981hybrid]. Together, these works established the Hamiltonian and Lagrangian foundations of modern acoustic and seismic wave propagation.
@@
Subsequent advances in semiclassical and asymptotic analysis, notably the Maslov canonical operator approach [maslov2001semi, kravtsov1990geometrical, kravtsov2012caustics, dobrokhotov2017new], have provided powerful mathematical tools for describing wavefields in weakly inhomogeneous and dispersive media. These techniques, based on the theory of canonical transformations in phase space, yield compact asymptotic representations of solutions near caustics and turning points, where traditional WKB or geometric acoustics approximations fail. Further refinements of this framework have been developed in works on operator separation of variables [belov2006operator, petrov2019application] and modern integral representations of the Maslov operator [dobrokhotov2017new], leading to robust methods for analyzing adiabatic and non-adiabatic mode evolution in three-dimensional shallow-water waveguides [petrov2019application].
@@
Parallel developments in optical and electromagnetic pulse theory have introduced analogous Hamiltonian structures for dispersive and lossy media. The analysis of group velocity dispersion, pulse-front tilt, and angular dispersion in ultrashort optical pulses [topp1975group, bor1985group, hebling1996derivation, hebling2008generation, osvay2004angular, kiselev2012nonparaxial] has strong mathematical correspondence with acoustic phase and amplitude evolution in horizontally varying environments. These analogies support the application of space–time ray methods and Hamiltonian formalisms beyond conservative systems, enabling cross-domain interpretations of wavefront dynamics in both acoustics and optics.
@@
Building on this theoretical foundation, we construct a non-ideal model for acoustic pulse propagation in a shallow-water waveguide that incorporates geometric and frequency dispersion, as well as partial energy transmission through the seabed. Starting from the general three-dimensional wave equation, we derive an $\varepsilon$-pseudodifferential operator ($\varepsilon$-PDO) formulation that separates fast (vertical) and slow (horizontal and temporal) variables. Applying the operator separation of variables method [belov2006operator] and the WKB–Maslov ansatz [maslov2001semi], we obtain eikonal and transport equations governing single-mode propagation within a Hamiltonian framework. This provides a physically consistent representation of mode dynamics and geometric spreading in non-ideal, weakly inhomogeneous environments.
@@
The use of non-self-adjoint and pseudo-Hermitian operators has become increasingly important in the analysis of dissipative and open wave systems. In the context of underwater acoustics, such operators naturally arise when accounting for partial reflection and energy leakage at the seabed [belov2006operator, petrov2019application]. Mathematically, this leads to the appearance of complex eigenvalues and biorthogonal mode structures [bender1998real, mostafazadeh2002pseudo, moiseyev2011nonhermitian], which can describe the attenuation and phase shift of propagating acoustic modes. The spectral properties of these non-self-adjoint systems differ fundamentally from those of Hermitian ones, as eigenfunctions are no longer orthogonal in the classical sense but form dual pairs associated with adjoint operators [berry2004physics, longhi2010pt, longhi2017parity].
@@
These theoretical developments have deep connections to semiclassical and Hamiltonian mechanics. The introduction of complex-valued Hamiltonians [berry2004physics, longhi2017parity] allows one to model the evolution of wavefronts in media with dissipation and gain while preserving the geometric structure of phase space. Similar formalisms have been successfully applied in optical and quantum contexts to describe PT-symmetric and lossy systems, providing valuable mathematical tools for the analysis of non-ideal acoustic waveguides. Integrating these ideas into the $\varepsilon$-PDO and WKB frameworks presented here enables a unified treatment of energy transport, geometric dispersion, and dissipation within a single Hamiltonian formalism.
@@
The resulting system provides explicit analytical and numerical expressions for acoustic fronts — including amplitude, phase, and time fields — under different bottom reflection conditions. This framework unifies and extends classical ray and mode theories, offering a consistent mathematical approach for describing non-ideal shallow-water acoustics and establishing a theoretical bridge between semiclassical analysis, dissipative wave mechanics, and ocean acoustics modeling.
@@ 1.4 $\varepsilon$-PDO and WKB-ansatz
Finally, we write simplified form of 1.45:
@@ 2.1 Real eigenvalues
For eigenproblem 2.1 we have, following general horizontal-vertical decomposition ideas [burridge2005horizontal, weinberg1974horizontal]:
@@ 2.2 Equation for WKB-ansatz
We now derive equations for amplitude and phase. We start from 2.6 and expand all objects with respect to small parameter up to the term $O(\varepsilon^{2})$
@@
Applying previous results for expansion 1.48, we have (omitting universal exponential factor)
@@ 4.1 Eikonal equation
where $\bm{r}(\sigma,\mu),\bm{p}(\sigma,\mu)$ are solutions of system 4.22 with initial values $\bm{r}_{0}(\mu),\bm{p}_{0}(\mu)$. Trajectories $\bm{f}(\sigma)$ in phase space we would call characteristics, while their projections $\bm{r}(\sigma)$ - rays. From the general theory we know that trajectories $\bm{f}(\cdot,\mu)$ and $\bm{f}(\cdot,\mu^{\prime})$ do not intersect, while corresponding rays can (and will, forming caustics). Pairs $\mathfrak{r}:=(\sigma,\mu)$ we would naturally call ray-centered coordinates or simply ray coordinates. For more detailed description we need to introduce notations for Jacobi matrices:
@@
Solution $\varphi$ of 4.1 for given characteristic $\bm{f}(\sigma,\mu)$ can be be found via formula:
@@ 4.2 Variation equations
Due to the fact, that Jacobi matrices have full rank (linearly independent columns), we can define Moore–Penrose (left) inverse matrices by the formula
@@ 7.2 Numerical example
8 Outlook
The present study extends the theoretical foundations of acoustic pulse propagation in shallow-water environments by incorporating non-self-adjoint operators and $\varepsilon$-pseudodifferential formulations. This approach captures essential non-ideal effects such as bottom interaction, mode leakage, and geometric dispersion—phenomena critical for realistic modeling of underwater acoustic fields.
@@ 8 Outlook
Second, numerical implementation and comparison with field data will be crucial to validate the theoretical predictions derived from the $\varepsilon$-PDO formalism. The development of hybrid models combining ray-based Hamiltonian dynamics with parabolic equation solvers or normal-mode codes (e.g., RAM and KRAKEN [collins1993splitstep, rogers2004ram]) could significantly improve computational efficiency and accuracy in coastal acoustics applications.
@@
Data Availibility
No datasets were generated or analyzed during the current study.
Conflict of interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.

### v2

@@ Abstract
Keywords: shallow-water acoustics, ray tracing, semiclassical analysis, non-self-adjoint operators, dispersion
@@ Acknowledgments
We dedicate this work to the memory of Prof. Vasilii Mikhailovich Babich (1930 – 2025), head of the St.Petersburg school in the theory of diffraction and wave propagation (PDMI RAS/LOMI and SPbU/LGU), whose foundational contributions shaped the modern asymptotic methods in wave theory [belishev2021vasilii, 11263469].
Aleksandr Kaplun is supported by the Center for Integration in Science, Israel Ministry of Aliyah and Integration and the post-doctoral scholarship at the Leon H. Charney School of Marine Sciences, University of Haifa. Boris Katsnelson is supported by the Israel Science Foundation, grant ISF-946/20.
1 Introduction
@@
Classical formulations of underwater sound propagation rely on ray and mode theories [weinberg1974horizontal, burridge2005horizontal, vcerveny2001seismic], in which the acoustic field is represented as a superposition of horizontally propagating rays and vertically standing modes. The theoretical foundation for such methods was established in the 1970s-1980s through the development of the space-time ray method by Babich, Molotkov, and Buldyrev [babich1998space, babich1980space], and the horizontal ray formalism of Weinberg and Burridge [weinberg1974horizontal]. These approaches were later generalized to complex and moving boundary problems [vcerveny1982space, kirpichnikova1983reflection, babich1984complex, farra1989ray, farra1993ray, connor1974complex] and further extended into hybrid ray-mode formulations for inhomogeneous ducts [felsen1981hybrid]. Together, these works established the Hamiltonian and Lagrangian foundations of modern acoustic and seismic wave propagation.
@@
Subsequent advances in semiclassical and asymptotic analysis, notably the Maslov canonical operator approach [maslov2001semi, kravtsov1990geometrical, kravtsov2012caustics, dobrokhotov2017new], have provided powerful mathematical tools for describing wavefields in weakly inhomogeneous and dispersive media. These techniques, based on the theory of canonical transformations in phase space, yield compact asymptotic representations of solutions near caustics and turning points, where traditional WKB or geometric acoustics approximations fail. Further refinements of this framework have been developed in works on operator separation of variables [belov2006operator] and modern integral representations of the Maslov operator [dobrokhotov2017new], leading to robust methods for analyzing adiabatic and non-adiabatic mode evolution in three-dimensional shallow-water waveguides [petrov2019application].
@@
Parallel developments in optical and electromagnetic pulse theory have introduced analogous Hamiltonian structures for dispersive and lossy media. The analysis of group velocity dispersion, pulse-front tilt, and angular dispersion in ultrashort optical pulses [topp1975group, bor1985group, hebling1996derivation, hebling2008generation, osvay2004angular, kiselev2012nonparaxial, katsnelson2018variability, katsnel2012space] has strong mathematical correspondence with acoustic phase and amplitude evolution in horizontally varying environments. These analogies support the application of space-time ray methods and Hamiltonian formalisms beyond conservative systems, enabling cross-domain interpretations of wavefront dynamics in both acoustics and optics.
@@
Building on this theoretical foundation, we construct a non-ideal model for acoustic pulse propagation in a shallow-water waveguide that incorporates geometric and frequency dispersion, as well as partial energy transmission through the seabed. Starting from the general three-dimensional wave equation, we derive an $\varepsilon$-pseudodifferential operator ($\varepsilon$-PDO) formulation that separates fast (vertical) and slow (horizontal and temporal) variables. Applying the operator separation of variables method [belov2006operator] and the WKB-Maslov ansatz [maslov2001semi], we obtain eikonal and transport equations governing single-mode propagation within a Hamiltonian framework. This provides a physically consistent representation of mode dynamics and geometric spreading in non-ideal, weakly inhomogeneous environments.
@@
The use of non-self-adjoint and pseudo-Hermitian operators has become increasingly important in the analysis of dissipative and open wave systems. In the context of underwater acoustics, such operators naturally arise when accounting for partial reflection and energy leakage at the seabed [belov2006operator, petrov2019application]. Mathematically, this leads to the appearance of complex eigenvalues and biorthogonal mode structures [bender1998real, mostafazadeh2002pseudo, moiseyev2011nonhermitian], which can describe the attenuation and phase shift of propagating acoustic modes. The spectral properties of these non-self-adjoint systems differ fundamentally from those of Hermitian ones, as eigenfunctions are no longer orthogonal in the classical sense but form dual pairs associated with adjoint operators [berry2004physics].
@@
These theoretical developments have deep connections to semiclassical and Hamiltonian mechanics. The introduction of complex-valued Hamiltonians [berry2004physics] allows one to model the evolution of wavefronts in media with dissipation and gain while preserving the geometric structure of phase space. Similar formalisms have been successfully applied in optical and quantum contexts to describe PT-symmetric and lossy systems, providing valuable mathematical tools for the analysis of non-ideal acoustic waveguides. Integrating these ideas into the $\varepsilon$-PDO and WKB frameworks presented here enables a unified treatment of energy transport, geometric dispersion, and dissipation within a single Hamiltonian formalism.
@@
The resulting system provides explicit analytical and numerical expressions for acoustic fronts – including amplitude, phase, and time fields – under different bottom reflection conditions. This framework unifies and extends classical ray and mode theories, offering a consistent mathematical approach for describing non-ideal shallow-water acoustics and establishing a theoretical bridge between semiclassical analysis, dissipative wave mechanics, and ocean acoustics modeling.
@@ 2.4 $\varepsilon$-PDO and WKB-ansatz
Finally, we write simplified form of 2.45:
@@ 3.1 Real eigenvalues
For eigenproblem 3.1 we have, following general horizontal-vertical decomposition ideas [burridge2005horizontal, weinberg1974horizontal]:
@@ 3.2 Equation for WKB-ansatz
We now derive equations for amplitude and phase. We start from 3.6 and expand all objects with respect to small parameter up to the term $O(\varepsilon^{2})$
@@
Applying previous results for expansion 2.48, we have (omitting universal exponential factor)
@@ 5.1 Eikonal equation
where $\bm{r}(\sigma,\mu),\bm{p}(\sigma,\mu)$ are solutions of system 5.22 with initial values $\bm{r}_{0}(\mu),\bm{p}_{0}(\mu)$. Trajectories $\bm{f}(\sigma)$ in phase space we would call characteristics, while their projections $\bm{r}(\sigma)$ - rays. From the general theory we know that trajectories $\bm{f}(\cdot,\mu)$ and $\bm{f}(\cdot,\mu^{\prime})$ do not intersect, while corresponding rays can (and will, forming caustics). Pairs $\mathfrak{r}:=(\sigma,\mu)$ we would naturally call ray-centered coordinates or simply ray coordinates. For more detailed description we need to introduce notations for Jacobi matrices:
@@
Solution $\varphi$ of 5.1 for given characteristic $\bm{f}(\sigma,\mu)$ can be be found via formula:
@@ 5.2 Variation equations
Due to the fact, that Jacobi matrices have full rank (linearly independent columns), we can define Moore-Penrose (left) inverse matrices by the formula
@@ 8.2 Numerical example
9 Conclusion
The present study extends the theoretical foundations of acoustic pulse propagation in shallow-water environments by incorporating non-self-adjoint operators and $\varepsilon$-pseudodifferential formulations. This approach captures essential non-ideal effects such as bottom interaction, mode leakage, and geometric dispersion?phenomena critical for realistic modeling of underwater acoustic fields.
@@ 9 Conclusion
Second, numerical implementation and comparison with field data will be crucial to validate the theoretical predictions derived from the $\varepsilon$-PDO formalism. The development of hybrid models combining ray-based Hamiltonian dynamics with parabolic equation solvers or normal-mode codes (e.g., KRAKEN [collins1993splitstep]) could significantly improve computational efficiency and accuracy in coastal acoustics applications.